2. Suppose IQ scores were obtained from randomly selected twins. For 20 such pairs of​ people, the linear correlation coefficient is 0.849 and the equation of the regression line is ŷ=0.68+0.99x​, where x represents the IQ score of the twin born first. ​Also, the 20 x values have a mean of 100.9 and the 20 y values have a mean of 100.5. What is the best predicted IQ of the twin born second​, given that the twin born first has an IQ of 105​? Use a significance level of 0.05.

The best predicted IQ of the twin born second is ____

​(Round to two decimal places as​ needed.)

3. A sample of 40 women is​ obtained, and their heights​ (in inches) and pulse rates​ (in beats per​ minute) are measured. The linear correlation coefficient is 0.284 and the equation of the regression line is ŷ=17.9+0.920x where x represents height. The mean of the 40 heights is 63.2 in and the mean of the 40 pulse rates is 73.3 beats per minute. Find the best predicted pulse rate of a woman who is 73 in tall. Use a significance level of a=0.05.

The best predicted pulse rate of a woman who is 73 in tall is ____ beats per minute.

​(Type an integer or decimal rounded to two decimal places as​ needed.)
Answer #1

Ans:

1)Given that

ŷ=0.68+0.99x

when x=105

The best predicted IQ of the twin born second is:

ŷ=0.68+0.99*105=104.63

2)

ŷ=17.9+0.920x

when x=73

The best predicted pulse rate of a woman who is 73 in tall is

ŷ=17.9+0.920*73 =85.06